# Discovering Interpretable Programmatic Policies via Multimodal LLM-assisted Evolutionary Search

### v1

@@ Abstract
Interpretability and high performance are essential goals in designing control policies, particularly for safety-critical tasks. Deep reinforcement learning has greatly enhanced performance, yet its inherent lack of interpretability often undermines trust and hinders real-world deployment. This work addresses these dual challenges by introducing a novel approach for programmatic policy discovery, called Multimodal Large Language Model-assisted Evolutionary Search (MLES). MLES utilizes multimodal large language models as policy generators, combining them with evolutionary mechanisms for automatic policy optimization. It integrates visual feedback-driven behavior analysis within the policy generation process to identify failure patterns and facilitate targeted improvements, enhancing the efficiency of policy discovery and producing adaptable, human-aligned policies. Experimental results show that MLES achieves policy discovery capabilities and efficiency comparable to Proximal Policy Optimization (PPO) across two control tasks, while offering transparent control logic and traceable design processes. This paradigm overcomes the limitations of predefined domain-specific languages, facilitates knowledge transfer and reuse, and is scalable across various control tasks. MLES shows promise as a leading approach for the next generation of interpretable control policy discovery.
@@ 1 Introduction
Interpretability and high performance are two central challenges in the design of policies for control tasks (Milani et al. 2024). In recent years, deep reinforcement learning (DRL) has achieved remarkable performance across various domains. However, two fundamental challenges persist, stemming from the neural network-based nature of DRL policies (Vouros 2022; Hickling et al. 2023; Puiutta and Veith 2020). First, such policies are opaque black boxes, offering limited transparency into their decision-making processes. This lack of interpretability undermines trust, verification, and adoption, particularly in safety-critical applications such as autonomous driving and healthcare (Perez-Cerrolaza et al. 2024). Second, the policy learning process in DRL relies on gradient-based optimization over high-dimensional parameter spaces, making it difficult to analyze, intervene in, or reuse the learned knowledge. These limitations have driven the ongoing pursuit of policies that are not only high-performing but also transparent, verifiable, and human-understandable.
@@
The recent rise of large language models (LLMs) offers a promising opportunity to develop policies that are both interpretable and high-performing. LLMs possess strong capabilities in contextual understanding, reasoning, and generation, particularly excelling in coding-related tasks (Achiam et al. 2023). Building on this foundation, the LLM-assisted Evolutionary Search (LES) paradigm has emerged as a powerful framework for automated design. LES combines the generative and reasoning capabilities of LLMs with the iterative optimization strengths of Evolutionary Computation (EC)(Liu et al. 2024a). In this paradigm, LLMs serve as mutation or crossover operators, guided by prompt templates and evaluation feedback, to iteratively refine candidate designs (Romera-Paredes et al. 2024; Liu et al. 2024a; Ye et al. 2024). LES has shown success in the automated discovery of code, algorithms, and heuristics. In the RL domain, recent studies such as Eureka et al. (Ma et al. 2024; Kwon et al. 2023; Sun et al. 2024; Masadome and Harada 2025) have demonstrated that LES can be used to evolve reward functions that improve agent performance.
Figure 1: Overview of methodological differences. (a) Standard RL: agents learn via reward-guided interaction with environments. (b) Existing LES (LLM-assisted Evolutionary Search) in the RL domain: agents are improved indirectly by reshaping reward functions using LLMs. (c) Our MLES: directly evolves policies by integrating behavior analysis during the policy discovery process.
@@
Inspired by these advances, we pose the question of whether LES can directly synthesize functional and interpretable programmatic policies, rather than merely designing auxiliary components like reward functions. This direction holds the potential to yield agents that are both effective and transparent, thus bridging the gap between high-performance black-box policies and human-understandable control logic. Building on this, we further investigate how LES can be enhanced to generate more reliable and generalizable policies, while also improving the efficiency of the policy discovery process.
@@
To address these questions, we propose Multimodal LLM-assisted Evolutionary Search (MLES), a novel framework for the automatic discovery of programmatic policies. MLES extends the LES paradigm in two key ways. First, it targets the direct evolution of policies, where policies are expressed as executable programs enriched with natural language rationales. These are developed through natural language interactions with LLMs, enabling the synthesis of policies that are both semantically meaningful and easily interpretable. Second, MLES introduces behavior analysis into the evolutionary loop. By grounding policy evaluation in richer signals beyond scalar quantitative metrics, MLES enables more precise, adaptable, and human-aligned policy evolution. This enhances the search efficiency and makes the policy discovery process more transparent and traceable. In summary, this paper makes the following contributions:
@@
(1) We propose the MLES framework, which integrates multimodal LLMs (MLLMs) with EC to directly synthesize interpretable programmatic policies through interaction with the environment. Unlike existing LES-based approaches that focus on reward shaping, MLES enables end-to-end policy discovery. An overview comparison is shown in Fig. 1.
@@
(2) We present a prototypical instantiation of MLES and evaluate it on two standard RL benchmarks: Lunar Lander and Car Racing. Experimental results show that MLES produces effective policies with transparent control logic and traceable design processes, achieving performance comparable to Proximal Policy Optimization (PPO).
@@
(3) We conduct extensive analyses of MLES in terms of effectiveness, search efficiency, and generalization, and investigate how different forms of behavioral evidence and prompt design influence the policy evolution process.
2 Related Work
2.1 Learning Interpretable Policies
Existing methods can be broadly categorized as either post hoc explanation and inherently interpretable modeling (Vouros 2022; Glanois et al. 2024). Post hoc methods attempt to explain trained black-box policies using human-understandable structures such as decision trees (Bastani, Pu, and Solar-Lezama 2018; Coppens et al. 2019; Gokhale et al. 2024; Kohler et al. 2024), mathematical formulas (Hein, Udluft, and Runkler 2018; Landajuela et al. 2021; Hazra and De Raedt 2023), finite-state machines (Koul, Greydanus, and Fern 2018; Inala et al. 2020), or domain-specific languages (DSLs) (Verma et al. 2018; Verma 2019). These methods typically rely on expert demonstrations or oracle supervision, using supervised or imitation learning to mimic the behavior (Verma et al. 2019; Cheng, Kolobov, and Agarwal 2020). Although achieving near-perfect performance, they offer limited insight into proactively constructing or improving policies, and tend to lack generalization beyond the oracle’s domain (Gu et al. 2025). Inherently interpretable approaches, by contrast, aim to discover interpretable policies directly from environment interactions. These policies are represented as formulas (Hein, Udluft, and Runkler 2019; Lyu et al. 2019), logic rules (Jiang and Luo 2019; Delfosse et al. 2023; Glanois et al. 2022), decision trees (Silva et al. 2020; Topin et al. 2021; Pace, Chan, and van der Schaar 2022; Wu et al. 2024), and DSLs (Trivedi et al. 2021; Qiu and Zhu 2022; Liu et al. 2023; Gu et al. 2024; Liu et al. 2024b). Although these methods improve transparency, they are often constrained by static grammars, handcrafted templates, and limited symbolic search spaces, which restrict their expressiveness, adaptability, and scalability in complex environments compared to DRL. Many further assume program sketches (Cao et al. 2022), reducing flexibility across diverse control tasks.
Our proposed MLES framework belongs to the latter category but substantially broadens its scope in three key ways. First, it represents policies as executable programs augmented with natural language explanations, enabling both transparency and semantic introspection. Second, it learns purely from environment interaction, without reliance on expert data or oracle guidance. Third, it replaces static grammars or handcrafted operators with LLMs, allowing for the dynamic generation of expressive and diverse control logic through linguistic reasoning. These features collectively enable the synthesis of policies that are human-readable and capable of expressing more precise decision logic.
2.2 LLM-assisted Evolutionary Search
EC has long served as a powerful paradigm for black-box optimization, particularly in non-differentiable or combinatorial domains (Bäck, Fogel, and Michalewicz 1997). Recent advances have integrated LLMs into the evolutionary loop, giving rise to a new class of frameworks known as LES. By employing LLMs as evolutionary operators, LES shifts the search space from numerical or symbolic encodings to linguistic representations, enabling the evolution of code or other structured artifacts. This paradigm has demonstrated success in domains such as code generation (Hemberg, Moskal, and O’Reilly 2024), scientific discovery (Romera-Paredes et al. 2024; Shojaee et al. 2024), and algorithm design (Liu et al. 2024a; Ye et al. 2024; Aglietti et al. 2024). In RL, LES has been primarily applied to the indirect optimization of agents by evolving reward functions (Ma et al. 2024; Hazra et al. 2024), thereby improving the performance of downstream neural policies. However, the resulting policies remain neural and opaque.
Figure 2: An overview of the MLES framework. The left side of the MLES illustrates the evolutionary search loop, while the right side details the structure and construction of an evolutionary individual. The red module on the far right exemplifies a method for generating behavioral evidence. During each search step, a subset of parent individuals is selected from a policy pool and used by the prompt sampler to create a multimodal few-shot prompt. MLLMs reason over this prompt to generate a new offspring policy. The offspring is then evaluated and visualized, resulting in the creation of a new individual that is added to the policy pool and managed accordingly. This iterative process continues, enabling the ongoing refinement of the policies.
In contrast, our work extends LES toward the direct synthesis of programmatic policies, shifting the focus from auxiliary reward function design to direct policy discovery from environment interaction. Furthermore, we incorporate behavior analysis into the evolutionary loop, allowing the search to be guided by richer information beyond scalar rewards. This enables a more informed and transparent form of policy discovery, in which both the resulting policy and its design rationale are accessible for inspection. To the best of our knowledge, this is the first known effort to unify LLMs, evolutionary search, and multimodal evaluation for the direct synthesis of interpretable policies in RL tasks.
2.3 Multimodal Large Language Models
MLLMs extend the capabilities of LLMs by integrating visual and textual modalities, enabling grounded understanding and vision-language reasoning (Caffagni et al. 2024). Recent models such as GPT-4V (Hurst et al. 2024) and Qwen2.5-VL (Bai et al. 2025) demonstrate strong performance on tasks requiring joint visual-language comprehension for planning, decision-making, and high-level control. These advances have led to increasing applications of MLLMs in planning and optimization. Some approaches employ MLLMs for real-time action selection and short-horizon planning in interactive environments (Zheng et al. 2024; Szot et al. 2025). Others leverage MLLMs as optimizers that iteratively refine solutions through visual feedback (Elhenawy et al. 2024a, b; Zhao and Cheong 2025), or use them for scene understanding in simulated and real-world domains (Li et al. 2025; Kambara et al. 2024; Hori et al. 2025). More recently, MLLMs have also been used to adapt reward functions based on visual cues (Narin 2024; Wang et al. 2025; Cuzin-Rambaud et al. 2025), thereby improving agent learning in complex RL settings.
These developments suggest that MLLMs are now capable of combining visual perception with contextual reasoning, a key capacity that aligns closely with our demands of interpretable policy discovery. In the MLES framework, we leverage the vision-language capabilities of MLLMs to guide the policy search process with rich information, including policy execution traces, in-depth analyses, and promising corrective suggestions. This approach facilitates a more transparent and traceable policy discovery, ultimately supporting the synthesis of reliable, human-aligned policies.
@@ 3 Methodology
This section presents the MLES framework, designed to automatically discover high-performing and interpretable programmatic policies for control tasks.
@@ 3.1 Problem Definition
Policy discovery refers to the process of searching for high-quality policies within a predefined policy space, with the goal of maximizing expected performance in a given environment. Formally, given a policy space $\Pi$ and an evaluation function $F(\pi)$ that measures the quality of a policy $\pi\in\Pi$, the policy discovery problem can be formulated as the following optimization problem:
@@
Different policy discovery methods achieve this by varying the policy space $\Pi$ and the optimization mechanisms used. For instance, DRL methods employ optimizers to search for optimal neural network parameters within a continuous parameter space. Genetic Programming combined with DSLs uses evolutionary algorithms to explore and identify the optimal combinations of predefined grammars within a discrete grammar space.
@@ 3.2 Framework Overview
The MLES framework orchestrates a closed-loop evolutionary process to iteratively search for programmatic policies within the policy space, as illustrated in Fig. 2. The framework requires the following configurable inputs to facilitate the discovery: (1) a selection of MLLMs for policy generation, (2) a formal task description, (3) an initial set of programmatic policies, (4) an evaluator for executing policies in the environment, and (5) a summarizer for generating interpretable behavioral evidence from policy executions. The framework consists of the following six core components:
Evolutionary search loop.
@@
The evolutionary search loop serves as the main control mechanism that orchestrates the entire evolutionary process. It manages iterations and maintains a dynamic balance between exploration and exploitation via configurable evolutionary operators, ensuring an effective search within the policy space.
Policy pool.
@@
The policy pool acts as a dynamic repository that holds the current population of policies. It is responsible for selecting appropriate parent candidates based on the evolutionary operator and integrating newly generated offspring into the population, thereby facilitating the continuous evolution of the policy set.
Evaluators.
@@
The evaluators are responsible for executing policies in the target environment. They output both quantitative performance metrics (e.g., episode rewards) and raw behavior traces. To improve efficiency and minimize delays in the policy search process, this evaluation procedure can be conducted in parallel.
Summarizer.
@@
The summarizer converts raw behavior traces collected by the evaluators into interpretable behavioral evidence. This evidence captures qualitative aspects of policy execution that go beyond scalar performance metrics, offering a deeper understanding of the policy’s behavior. This component is crucial for enabling informed policy modifications by the MLLMs and represents a key innovation of MLES.
Prompt sampler.
@@
The prompt sampler constructs multimodal few-shot prompts tailored to the current evolutionary operator by integrating selected parent policies and their corresponding interpretable behavioral evidence. These prompts provide essential context and guidance, directing the MLLMs to propose meaningful policies aligned with the intended transformation, thus steering the evolution toward promising directions.
LLMs ensemble.
@@
This component consists of a configurable set of MLLMs. These MLLMs reason over the few-shot prompts, leveraging their advanced vision-language understanding, reasoning, and code generation capabilities to generate new candidate policies.
@@
Through the iterative refinement process, MLES progressively discovers a diverse set of interpretable and high-performing programmatic policies. Concurrently, this process explores a wide range of both successful and suboptimal policies, yielding valuable insights that enhance human understanding of the target task and provide a foundation for knowledge transfer and reuse in related control scenarios.
3.3 Policy Representation
@@
During the search process, each candidate policy is treated as an individual, structured as a tuple consisting of four elements that capture the policy’s logic, intent, performance, and behavioral characteristics: Code, Thought, Quantitative Metrics, and Interpretable Behavioral Evidence (IBE).
Code.
@@
Each policy is implemented as an executable Python program that defines the agent’s decision-making logic. This representation supports automated evaluation within the evolutionary loop and ensures direct compatibility with the environment. Moreover, expressing policies in code offers high transparency and modularity, making them human-readable, easy to debug, and conducive to reuse or modification during the search process.
Thought.
@@
The thought is a concise natural language summary that captures the underlying rationale or design intent of a policy, highlighting key strategies that guide the agent’s behavior. By providing a high-level semantic abstraction, thoughts help LLMs interpret the associated code more effectively and support deeper reasoning during policy generation. Prior work (Liu et al. 2024a) has shown that incorporating such descriptions into few-shot prompts can significantly improve the efficiency and effectiveness of LES.
Quantitative metrics.
@@
Quantitative metrics are task-specific numerical evaluations obtained by executing the policy in the environment. These metrics serve as direct indicators of policy’s performance (i.e., $F(\cdot)$ in 1), such as episode rewards, success rates, or other user-defined scores tailored to the tasks. They provide a consistent basis for ranking and selecting candidate policies, ensuring high-performing individuals are effectively identified and prioritized for further refinement.
IBE.
@@
While quantitative metrics are essential for measuring performance, they provide only a partial understanding of a policy’s behavior pattern and often fail to offer actionable guidance for improvement. To address this limitation, MLES introduces IBE, which captures the overall behavior of a policy or highlights key events that help identify areas for improvement. By analyzing IBE, MLLMs can observe policy outcomes in detail, accurately identifying failure modes such as incorrect decisions or reward hacking (e.g., exploiting bugs for undeserved rewards), thus enabling more targeted policy modifications and corrections. This provides richer information for MLLMs to reason over, facilitating the discovery of human-aligned policies. This design mirrors how human experts refine policies: they do not merely rely on evaluating numerical scores but also examine behavioral patterns (e.g., failure modes, unexpected interactions) to identify areas for improvement. The form of IBE is flexible, including images, videos, and even text-based state sequences, and more. Examples of these formats and more detailed descriptions can be found in Appendix C.

### v2

# Multimodal LLM-assisted Evolutionary Search for Programmatic Control Policies

@@ Abstract
Deep reinforcement learning has achieved impressive success in control tasks. However, its policies, represented as opaque neural networks, are often difficult for humans to understand, verify, and debug, which undermines trust and hinders real-world deployment. This work addresses this challenge by introducing a novel approach for programmatic control policy discovery, called Multimodal Large Language Model-assisted Evolutionary Search (MLES). MLES utilizes multimodal large language models as programmatic policy generators, combining them with evolutionary search to automate policy generation. It integrates visual feedback-driven behavior analysis within the policy generation process to identify failure patterns and guide targeted improvements, thereby enhancing policy discovery efficiency and producing adaptable, human-aligned policies. Experimental results demonstrate that MLES achieves performance comparable to Proximal Policy Optimization (PPO) across two standard control tasks while providing transparent control logic and traceable design processes. This approach also overcomes the limitations of predefined domain-specific languages, facilitates knowledge transfer and reuse, and is scalable across various tasks, showing promise as a new paradigm for developing transparent and verifiable control policies.
@@ 1 Introduction
High performance and transparency are two central challenges in the design of policies for control tasks (Milani et al., 2024). In recent years, deep reinforcement learning (DRL) has achieved remarkable performance across various domains. However, two fundamental challenges persist, stemming from the opaque, neural network-based nature of DRL policies (Vouros, 2022; Hickling et al., 2023; Puiutta & Veith, 2020). First, such policies function as black boxes, offering limited insight into their decision-making processes. This lack of transparency undermines trust, complicates verification, and hinders adoption, particularly in safety-critical applications such as autonomous driving and healthcare (Perez-Cerrolaza et al., 2024). Second, the policy learning process in DRL relies on gradient-based optimization over high-dimensional parameter spaces, making it difficult to analyze, intervene in, or reuse the learned knowledge. These limitations have driven the ongoing pursuit of policies that are not only high-performing but also transparent, verifiable, and human-readable.
@@
The recent rise of large language models (LLMs) offers a promising opportunity to develop control policies that are both transparent and high-performing. LLMs possess strong capabilities in contextual understanding, reasoning, and generation, particularly excelling in coding-related tasks (Achiam et al., 2023). Building on this foundation, the LLM-assisted Evolutionary Search (LES) paradigm has emerged as a powerful framework for automated design. LES combines the generative and reasoning capabilities of LLMs with the iterative optimization strengths of Evolutionary Computation (EC)(Liu et al., 2023a). In this paradigm, LLMs serve as mutation or crossover operators, guided by prompt templates and evaluation feedback, to iteratively refine candidate designs (Romera-Paredes et al., 2024; Liu et al., 2024a; Ye et al., 2024). LES has shown success in the automated discovery of code, algorithms, and heuristics. In the reinforcement learning domain, recent studies such as Eureka et al. (Ma et al., 2024; Kwon et al., 2023; Sun et al., 2024; Masadome & Harada, 2025) have demonstrated that LES can be used to shape reward functions that improve agent performance.
@@
Inspired by these advances, we pose the question of (1) whether LES can directly synthesize high-performance programmatic policies, rather than merely designing auxiliary components like reward functions. This direction holds the potential to yield agents that are both effective and transparent, thus bridging the gap between high-performance black-box models and human-understandable control logic. Building on this, we further investigate (2) how this process can be enhanced to generate more reliable policies, while also improving the efficiency of the policy discovery.
Figure 1: Overview of methodological differences. (a) Standard DRL: agents learn via reward-guided interaction with environments. (b) Our MLES: directly evolves programmatic policies by integrating behavior analysis during the EC-based policy discovery process.
@@
To address these questions, we propose Multimodal LLM-assisted Evolutionary Search (MLES), a novel framework for the automatic discovery of programmatic control policies. MLES extends the LES paradigm in two key ways. First, it targets the direct evolution of policies, where policies are expressed as executable programs enriched with natural language rationales. These are developed through natural language interactions with LLMs, enabling the synthesis of policies that are both semantically meaningful and easily understandable. Second, MLES introduces behavior analysis into the evolutionary loop. This enhances the search efficiency and makes the policy discovery process more transparent and traceable. By grounding policy evaluation in richer signals beyond scalar quantitative metrics, MLES enables more adaptable and human-aligned policy evolution. In summary, this paper makes the following contributions:
@@
(1) We propose the MLES framework, which integrates multimodal LLMs (MLLMs) with EC to directly synthesize programmatic control policies through interaction with the environment. Unlike existing LES-based approaches that focus on reward function shaping, MLES enables end-to-end policy discovery, akin to DRL methods. An overview comparison is shown in Fig. 1.
@@
(2) We present a prototypical instantiation of MLES and evaluate it on two standard RL benchmarks: Lunar Lander and Car Racing. Experimental results show that MLES produces effective policies with transparent control logic and traceable design processes, achieving performance comparable to a strong DRL baseline, Proximal Policy Optimization (PPO).
@@
(3) We conduct an extensive analysis of MLES regarding its effectiveness and search efficiency, complemented by thorough ablation studies. We also investigate how different forms of behavioral evidence and prompt designs influence the policy evolution process.
@@ 2.1 Problem Definition
Control policy discovery refers to the process of searching for high-quality policies within a predefined policy space, with the goal of maximizing expected performance in a given environment. Formally, given a policy space $\Pi$ and an evaluation function $F(\pi)$ that measures the quality of a policy $\pi\in\Pi$, the policy discovery problem can be formulated as the following optimization problem:
@@
Different policy discovery methods achieve this by varying the policy space $\Pi$ and the optimization mechanisms used. For instance, DRL methods employ optimizers to search for optimal neural network parameters within a continuous parameter space. Genetic Programming, combined with domain-specific languages (DSLs), uses evolutionary algorithms to explore and identify the optimal combinations of predefined grammars within a discrete grammar space.
@@ 2.2 Framework Overview
Figure 2: An overview of the MLES framework. The left side of the MLES illustrates the evolutionary search loop, while the right side details the structure and construction of an evolutionary individual. The red module on the far right exemplifies a method for generating behavioral evidence. During each search step, a subset of parent individuals is selected from a policy pool and used by the prompt sampler to create a multimodal few-shot prompt. MLLMs reason over this prompt to generate a new offspring policy. The offspring is then evaluated and visualized, resulting in the creation of a new individual that is added to the policy pool and managed accordingly.
@@
The MLES framework orchestrates a closed-loop evolutionary process to iteratively search for programmatic policies within the policy space, as illustrated in Fig. 2. The framework requires the following configurable inputs to facilitate the discovery: (1) a selection of MLLMs for policy generation, (2) a formal task description, (3) an initial set of programmatic policies, (4) an evaluator for executing policies in the environment, and (5) a summarizer for generating behavioral evidence from policy executions. The framework consists of the following six core components:
@@
Evolutionary search loop. The evolutionary search loop serves as the main control mechanism that orchestrates the entire evolutionary process. It manages iterations and maintains a dynamic balance between exploration and exploitation via configurable evolutionary operators, ensuring an effective search within the policy space.
@@
Policy pool. It acts as a dynamic repository that holds the current policy population. It is responsible for selecting appropriate parent policies based on the evolutionary operator and integrating newly generated offspring into the pool, thereby facilitating the continuous evolution of the policies.
@@
Evaluators. The evaluators are responsible for executing policies in the target environment. They output both quantitative performance metrics (e.g., episode rewards) and raw behavior traces. To improve efficiency and minimize delays, the policy evaluation procedure is often parallelized.
@@
Summarizer. It converts raw behavior traces collected by the evaluators into behavioral evidence. This evidence captures qualitative aspects of policy execution that go beyond scalar performance metrics, offering a deeper understanding of the policy’s behavior. This component is crucial for enabling informed policy modifications by the MLLMs and represents a key innovation of our MLES.
@@
Prompt sampler. It constructs multimodal few-shot prompts tailored to the current evolutionary operator by integrating selected parent policies and their corresponding behavioral evidence. These prompts provide essential context and guidance, directing the MLLMs to propose meaningful policies aligned with the intended transformation, thus steering the evolution toward promising directions.
@@
LLMs ensemble. This component consists of a configurable set of MLLMs. These MLLMs reason over the multimodal few-shot prompts, leveraging their advanced vision-language understanding, reasoning, and code generation capabilities to generate new candidate policies.
@@
Through the iterative refinement process, MLES progressively discovers a diverse set of high-performing programmatic policies. Concurrently, this process explores a wide range of both successful and suboptimal policies, yielding valuable insights that enhance human understanding of the target task and provide a foundation for knowledge transfer and reuse in related control scenarios.
2.3 Policy Representation in Evolutionary Search Process
@@
During the search process, each candidate policy is treated as an individual, structured as a tuple consisting of four elements that capture the policy’s logic, intent, performance, and behavioral characteristics: Code, Thought, Quantitative Metrics, and Behavioral Evidence.
@@
Code. Each policy is implemented as an executable Python program that defines the agent’s decision-making logic. This representation supports automated evaluation within the evolutionary loop and ensures direct compatibility with the environment. Moreover, expressing policies in code offers high transparency and modularity, making them human-readable, easy to debug, and conducive to reuse or modification during the search process.
@@
Thought. The thought is a concise natural language summary that captures the underlying rationale or design intent of a policy, highlighting key strategies that guide the agent’s behavior. By providing a high-level semantic abstraction, thoughts help LLMs interpret the associated code more effectively and support deeper reasoning during policy generation. Prior work (Liu et al., 2024a) has shown that incorporating such descriptions into prompts can improve the efficiency and effectiveness of LES.
@@
Quantitative metrics. Quantitative metrics are task-specific numerical evaluations obtained by executing the policy in the environment. These metrics serve as direct indicators of policy’s performance (i.e., $F(\cdot)$ in 1), such as episode rewards, success rates, or other user-defined scores tailored to the tasks. They provide a consistent basis for ranking and selecting candidate policies, ensuring high-performing individuals are effectively identified and prioritized for further refinement.
@@
Behavioral Evidence. (BE) While quantitative metrics are essential for measuring performance, they provide only a partial understanding of a policy and often fail to offer actionable guidance for policy improvement. To address this limitation, MLES introduces BE, which captures the overall behavior of a policy or highlights key events that help identify areas for improvement. By analyzing BE, MLLMs can observe policy outcomes in detail, accurately identifying failure modes such as incorrect decisions or reward hacking (e.g., exploiting bugs for undeserved rewards), thus enabling more targeted policy modifications and corrections. This provides richer information for MLLMs to work with, facilitating the discovery of human-aligned policies. This design is similar to how human experts refine policies: they do not merely rely on evaluating numerical scores but also examine behavioral patterns (e.g., failure modes, unexpected interactions) to identify areas for improvement. The format of BE is flexible, including images, videos, and even text-based state sequences, and more. Further examples illustrating the format and necessity of BE can be found in Appendix D.